solve the intergrating factt ;tdt+xdx=a2(tdx-xdx)/t^2+x^2
1
Expert's answer
2021-06-17T12:45:06-0400

Solve the integrating factor:

"tdt+xdx=\frac{a^2(tdx-xdt)}{t^2+x^2}"

Solution:

"(x-\frac{a^2t}{t^2+x^2})dx+(t+\frac{a^2x}{t^2+x^2})dt=0"

"Pdx+Qdt=0"

where "P=(x-\frac{a^2t}{t^2+x^2})" and "Q=(t+\frac{a^2x}{t^2+x^2})" .

Let's find "\frac{\partial P}{\partial t}" and "\frac{\partial Q}{\partial x}":

"\frac{\partial P}{\partial t}=-a^2\cdot\frac{x^2+t^2-2t^2}{(t^2+x^2)^2}=a^2\cdot\frac{t^2-x^2}{(t^2+x^2)^2}";

"\frac{\partial Q}{\partial x}=a^2\cdot\frac{t^2+x^2-2x^2}{(t^2+x^2)^2}=a^2\cdot\frac{t^2-x^2}{(t^2+x^2)^2}" .

Since "\frac{\partial P}{\partial t}=\frac{\partial Q}{\partial x}" than we can conclude that it is a total differential equation.

Its solution can be find as:

"U(x,t)=\int_{x_0}^x P(\tau,t_0)d\tau+\int_{t_0}^t Q(x,\tau)d\tau"

where "x_0" and "t_0" are any values that can be freely chosen. Let's choose "x_0=1", "t_0=0" .

"U(x,t)=\int_{1}^x (\tau-\frac{a^2\cdot0}{0^2+\tau^2})d\tau+\int_{0}^t (\tau+\frac{a^2x}{\tau^2+x^2})d\tau="

"\int_{1}^x \tau d\tau+\int_{0}^t (\tau+\frac{a^2x}{\tau^2+x^2})d\tau=" "\frac{x^2}{2}-\frac12+\frac{t^2}{2}+a^2\arctan{\frac {t}{x}}" .

So the general integral of the equation has the form:

"\frac{x^2}{2}+\frac{t^2}{2}+a^2\arctan{\frac {t}{x}}=const" .

Answer:

This equation is a total differential equation therefore, its solution can be found without an integrating factor or any constant value can be used as integrating factor in this case.

Solution: "\frac{x^2}{2}+\frac{t^2}{2}+a^2\arctan{\frac {t}{x}}=const" .